OpenStudy (anonymous):

Find the range if y = 3x + 12 for x = {0.4, 0.6, 1.8, 2.2}.

OpenStudy (nurali):

y = 3*(0.4) + 12 = 13.2 y = 3*(0.6) + 12 = 13.8 y = 3*(1.8) + 12 = 17.4 y = 3*(2.2) + 12 = 18.6 Range = {13.2, 13.8, 17.4, 18.6}

OpenStudy (anonymous):

Thanks

OpenStudy (nurali):

My Pleasure.

OpenStudy (nurali):

Welcome to Openstudy.